# REBEL: Reinforcement Learning via Regressing Relative Rewards

### v1

@@ 5.1 Natural Language Generation
We evaluate each method by its balance between reward model score and KL-divergence with the reference policy, testing the effectiveness of the algorithm in optimizing the regularized RL object. To evaluate the quality of the generation, we compute the winrate (Rafailov et al., 2023) against human references using GPT4888Specific API checkpoint used throughout this section: gpt-4-0613 (OpenAI, 2023). The winrate is computed from a randomly sampled subset ($10\%$) of the test set with a total of 600 samples. The prompt used to query GPT4 as well as an example response is shown in Appendix D.3.
5.1.1 Quality Analysis
@@
1.4B SFT 24.5% -0.52 -
DPO 43.8% 0.11 30.9
PPO 51.6% 1.73 29.1
REBEL 55.3% 1.87 32.4
2.8B SFT 28.4% -0.40 -
DPO 53.5% 2.41 66.5
PPO 67.2% 2.37 27.4
REBEL 70.3% 2.44 29.2
Table 1: Results on TL;DR Summarization for SFT, PPO, DPO, and REBEL using three metrics. The RM Score is computed using the reward model with the respective size and the winrate is evaluated by GPT4. The models are trained with low-rank adapters. The best-performing method for each size and metric is highlighted in bold and the second best is underlined. We note that REBEL outperforms all baselines here in terms of the winrate
6.9B
SFT DPO REINFORCE PPO RLOO ($k=2$) RLOO ($k=4$) REBEL
Winrate ($\uparrow$) 44.6% 68.2% 70.7% 77.6% 74.2% 77.9% 78.0%
@@
Table 2: Results on TL;DR Summarization on 6.9B models. We perform full-parameter training for all models. The best-performing method is highlighted in bold and the second best is underlined.
Figure 2: Plot of Reward vs KL-Divergence for 2.8B REBEL and PPO. We evaluate the models across the entire test set every 100 steps for 2,000 steps. Left: each point represents the average reward score and KL-divergence for a specific time step; the eclipse represents the confidence interval with 2 standard deviations. Right: we divide the KL distribution at the 2,000-step into $10$ bins with equal size and average the corresponding RM scores in each bin.
[…]
The trade-off between the reward model score and KL-divergence is shown in Figure 2. We evaluate the 2.8B REBEL and PPO every 400 gradient updates during training for 8,000 updates. The sample complexity of each update is held constant across both algorithms for fair comparison. For the left plot, each point represents the average divergence and score over the entire test set, and the eclipse represents the confidence interval with 2 standard deviations. As observed previously, PPO exhibits lower divergence, whereas REBEL shows higher divergence but is capable of achieving larger RM scores. Notably, towards the end of the training (going to the right part of the plot), REBEL and PPO have similar KL and RM scores. For the right plot in Figure 2, we analyze a single checkpoint for each algorithm at the end of training. For each algorithm, we group every generation from the test set by its KL distribution into $10$ equally sized bins and calculate the average of the corresponding RM score for each bin. We can see that REBEL achieves higher RM scores for generations with small divergence while requiring larger divergence for generations with the highest scores.
5.1.2 Runtime & Memory Analysis
@@
We analyze the runtime and peak memory usage for 2.8B models using PPO, DPO, RLOO, and REBEL. The runtime includes both the generation time and the time required for policy updates. Both runtime and peak memory usage are measured on A6000 GPUs using the same hyperparameters detailed in Appendix D.2. The methods in the plots are arranged in ascending order based on winrates. To the right of the dashed line, PPO, RLOO ($k=4$), and REBEL have the highest winrates, which are comparable among them.
@@
While DPO and REINFORCE require less time and memory, their performance does not match up to REBEL, as discussed in Section 5.1.1. RLOO ($k=2$) has similar runtime and memory usage as REBEL since we set $\mu=\pi_{t}$, making REBEL also generate twice per prompt. However, RLOO ($k=2$) has worse performance than REBEL. Compared to PPO and RLOO ($k=4$), REBEL demonstrates shorter runtimes and lower peak memory usage. PPO is slow and requires more memory because it needs to update both two networks: policy network and value network. RLOO ($k=4$) requires generating 4 responses per prompt which makes it slow and less memory efficient. Compared to the two baselines PPO and RLOO ($k=4$) that achieve similar winrates as REBEL, we see that REBEL is more computationally tractable. REBEL is also noticeably simpler to implement than PPO since it does not learn value networks or compute the advantage estimation.
Figure 3: Plot of runtime and memory usage for DPO, REINFORCE, RLOO, PPO, and REBEL. The runtime includes both time for generation and policy update for each batch. Runtime and memory usage are measured on A6000 GPUs. Baselines on the left-hand side of the dashed line have lower winrates. Methods on the right-hand side of the dashed line have similar winrates to REBEL, but REBEL is noticeably more computationally tractable and memory efficient than PPO and RLOO ($k=4$).
@@ 5.2 Image Generation
We also consider the setting of image generation, where, given a consistency model (Song et al., 2023a) and a target reward function, we seek to train the consistency model to output images which garner a higher reward. Specifically, we compare REBEL and PPO under the RLCM framework (Oertell et al., 2024).
Baselines:
@@
We compare REBEL to a clipped, policy gradient objective (Black et al., 2023; Fan et al., 2024; Oertell et al., 2024) with the aim to optimize aesthetic quality to obtain high reward from the LAION aesthetic score predictor (Schuhmann, 2022). This baseline does not use critics or GAE for advantage estimates. However, the clipping objective is clearly motivated by PPO, and thus, we simply name this baseline as PPO in this section.
Dataset:
We use 45 common animals as generation prompts similar to Black et al. (2023); Oertell et al. (2024) [^9] Dataset available at https://github.com/Owen-Oertell/rlcm.
Models:
We use the latent consistency model (Luo et al., 2023) distillation of the Dreamshaper v7 model [^10] Huggingface model card: SimianLuo/LCM_Dreamshaper_v7, a finetune of stable diffusion (Rombach et al., 2021).
Figure 4: Learning curves as a function of reward queries to the LAION aesthetic predictor. We report inter-quartile means (IQM) with 95% confidence intervals (CIs) across three seeds for both REBEL and PPO. The CIs were calculated with percentile bootstrap with stratified sampling over three random seeds.
Evaluation:
@@
We evaluate PPO and REBEL on its reward under the LAION aesthetic reward model for an equal number of reward queries/samples generated and an equal number of gradient updates. The aesthetic predictor is trained to predict human-labeled scores of images on a scale of 1 to 10. Images that tend to have the highest reward are artwork. Following the recommendations of Agarwal et al. (2021b), we report the inter-quartile mean with 95% confidence intervals for our reported results across three random seeds.
5.3 Quality Analysis
@@
Figure 5: Generated images using PPO and REBEL during an intermediate checkpoint. We note that at the same number of epochs, REBEL observes a higher reward under the reward model. This can further be seen by the more diverse background of images generated from REBEL with less training time.
@@
Figure 4 shows REBEL optimizes the consistency model faster during the beginning of training but eventually achieves similar performance to that of PPO. For our experiments, we tuned both batch size and learning rate for our algorithms, testing batch sizes of $[4,8,16]$ per gpu and learning rates $[1\text{e}-4,3\text{e}-4,6\text{e}-4,1\text{e}-3]$. Note, the main difference in implementation between PPO and REBEL is the replacement of the clipped PPO objective with our regression objective.
Qualitatively, we observe that eventually, both PPO and REBEL start to generate good-looking images but ignore the text prompt entirely. However, from just optimizing the reward function perspective, this behavior is not surprising since the objective does not encourage the maintenance of the consistency between the text prompt and the generated image. To maximize LAION-predicted aesthetic quality, both REBEL and PPO transform a model that produces plain images into one that produces artistic drawings. We found across multiple seeds that REBEL produced lush backgrounds when compared to PPO’s generations. Please see Appendix E.2 for more examples of generated images.
@@ 6 Related Work
Policy Gradients. Policy gradient (PG) methods (Nemirovskij and Yudin, 1983; Williams, 1992; Sutton et al., 1999; Konda and Tsitsiklis, 1999; Kakade, 2001; Schulman et al., 2017) are a prominent class of RL algorithms due to their direct, gradient-based policy optimization, robustness to model mis-specification (Agarwal et al., 2020), and scalability to modern AI applications from fine-tuning LLMs (Stiennon et al., 2022) to optimizing text-to-image generators (Oertell et al., 2024).
Broadly speaking, we can taxonomize PG methods into two families. The first family is based on REINFORCE (Williams, 1992) and often includes variance reduction techniques (Kool et al., 2019; Richter et al., 2020; Zhu et al., 2023). While prior work by Ahmadian et al. (2024) has shown that REINFORCE-based approaches can outperform more complex RL algorithms like PPO on LLM fine-tuning tasks like TL;DR, we find that a properly optimized version of PPO still out-performs a REINFORCE baseline. The second family is adaptive PG techniques that precondition the policy gradient (usually with the inverse of the Fisher Information Matrix) to ensure it is covariant to re-parameterizations of the policy, which include NPG (Kakade, 2001; Bagnell and Schneider, 2003) and its practical approximations like TRPO (Schulman et al., 2015a) and PPO (Schulman et al., 2017). Intuitively, the preconditioning ensures that we make small changes in terms of action distributions, rather than in terms of the actual policy parameters, leading to faster and more stable convergence. Unfortunately, computing and then inverting the Fisher Information Matrix is computationally intensive and therefore we often resort to approximations in practice, as done in TRPO. However, these approximations are still difficult to apply to large-scale generative models, necessitating even coarser approximations like PPO. In contrast, REBEL does not need any such approximations to be implemented at scale, giving us a much closer connection between theory and practice.
Reward Regression. The heart of REBEL is a novel reduction from RL to iterative squared loss regression. While using regression to fit either the reward (Peters and Schaal, 2007) or the value (Peng et al., 2019) targets which are then used to extract a policy have previously been explored, our method instead takes a page from DPO (Rafailov et al., 2023) to implicitly parameterize the reward regressor in terms of the policy. This collapses the two stage procedure of prior methods into a single regression step.
@@
The third family of PFT techniques combines elements from the previous two: it involves running an offline algorithm iteratively, collecting on-policy preference feedback from either a supervisor model (Rosset et al., 2024; Xiong et al., 2024; Guo et al., 2024) or from a preference model fit on human data (Calandriello et al., 2024). All of these approaches can be considered instantiations of the general SPO reduction proposed by Swamy et al. (2024), which itself can be thought of as a preference-based variant of DAgger (Ross et al., 2011). Recent work by Tajwar et al. (2024) confirms the empirical strength of these techniques. Our approach fits best into this family of techniques – we also iteratively update our model by solving a sequence of supervised learning problems over on-policy datasets. However, REBEL comes with several key differentiating factors from the prior work. First, we can run REBEL with datasets consisting of a mixture of on-policy and off-policy data with strong guarantees, enabling hybrid training, as previously explored in the RL (Song et al., 2023b; Ball et al., 2023; Zhou et al., 2023) and inverse RL (Ren et al., 2024) literature. Second, unlike all of the aforementioned works that regularize to the initial policy $\pi_{0}$ during updates, we perform conservative updates by regularizing $\pi_{t+1}$ to $\pi_{t}$. Thus, for the prior work, it is difficult to prove convergence or monotonic improvement as the current policy can just bounce around a ball centered at $\pi_{0}$, a well-known issue in the theory of approximate policy iteration (Kakade and Langford, 2002; Munos, 2003). In contrast, by incorporating the prior policy’s probabilities into our regression problem, we are able to prove stronger guarantees for REBEL.
7 Summary and Future Work

### v2

@@ 5.1 Summarization
Evaluation. We evaluate each method by its balance between reward model score and KL-divergence with the SFT policy, testing the effectiveness of the algorithm in optimizing the regularized RL objective. To evaluate the quality of the generation, we compute the winrate (Rafailov et al., 2023) against human references using GPT4 (OpenAI, 2023). The winrate is computed from a randomly sampled subset ($10\%$) of the test set with 600 samples. We report the average results over three seeds.
Figure 2: Plot of runtime and memory usage. Baselines on the left-hand side of the dashed line have lower winrates. Methods on the right-hand side of the dashed line have similar winrates to REBEL.
@@
1.4B SFT 24.9 (±$\pm 2.73$) -0.51 (±$\pm 0.05$) -
DPO 42.7 (±$\pm 1.79$) 0.10 (±$\pm 0.02$) 29.6 (±$\pm 0.63$)
Iterative DPO 47.2 (±$\pm 1.34$) 1.73 (±$\pm 0.05$) 29.7 (±$\pm 0.57$)
PPO 51.7 (±$\pm 1.42$) 1.74 (±$\pm 0.04$) 29.3 (±$\pm 0.61$)
REBEL 55.1 (±$\pm 1.35$) 1.84 (±$\pm 0.04$) 32.6 (±$\pm 0.59$)
2.8B SFT 28.2 (±$\pm 2.31$) -0.38 (±$\pm 0.06$) -
DPO 53.7 (±$\pm 1.63$) 2.40 (±$\pm 0.02$) 64.3 (±$\pm 1.25$)
Iterative DPO 63.1 (±$\pm 1.41$) 2.37 (±$\pm 0.03$) 28.1 (±$\pm 0.51$)
PPO 67.4 (±$\pm 1.30$) 2.37 (±$\pm 0.03$) 27.2 (±$\pm 0.55$)
REBEL 70.2 (±$\pm 1.32$) 2.44 (±$\pm 0.02$) 29.0 (±$\pm 0.60$)
Model size Algorithm Winrate ($\uparrow$)
6.9B SFT 45.2 (±$\pm 2.49$)
DPO 68.4 (±$\pm 2.01$)
REINFORCE 70.7
PPO 77.6
RLOO ($k=2$) 74.2
RLOO ($k=4$) 77.9
REBEL 78.1 (±$\pm 1.74$)
@@
Table 1: Results on TL;DR Summarization. Results are averaged over three seed and the standard deviations across seeds are in parentheses. The best-performing method for each size and metric is highlighted in bold and the second best is underlined. REBEL outperforms all baselines on winrate.
Quality Analysis. Table 1 presents a comparison between REBEL and baseline methods. Notably, REBEL outperforms all the baselines on RM score with 1.4B and 2.8B parameters with a slightly larger KL than PPO. In addition, REBEL achieves the highest winrate under GPT4 when evaluated against human references, indicating the benefit of regressing the relative rewards. An ablation analysis on parameter $\eta$ is in Appendix H and the trade-off between the reward model score and KL-divergence is discussed in Appendix I.
@@
Runtime & Memory Analysis. We analyze the runtime and peak memory usage for 2.8B models with REINFORCE, RLOO, PPO, DPO, Iterative DPO, and REBEL. The runtime includes both the generation time and the time required for policy updates. Both runtime and peak memory usage are measured on A6000 GPUs using the same hyperparameters detailed in Appendix F.1.5 for a batch of $512$ prompts. The measurements are averaged over $100$ batches. Methods are ascendingly ordered by winrate. To the right of the dashed line, PPO, RLOO ($k=4$), and REBEL have the highest winrates, which are comparable among them.
@@
While DPO and REINFORCE are more time and memory-efficient, their performance does not match up to REBEL, as shown in Table 1. RLOO ($k=2$) and Iterative DPO have similar runtime and memory usage as REBEL since we set $\mu=\pi_{t}$, making REBEL also generate twice per prompt. However, both methods have worse performance than REBEL. Compared to PPO and RLOO ($k=4$), REBEL demonstrates shorter runtimes and lower peak memory usage. PPO is slow and requires more memory since it needs to update two networks (the policy network and the value network). RLOO ($k=4$) requires four generations per prompt which makes it slow and less memory efficient. In summary, compared to the two baselines (PPO and RLOO ($k=4$)) that achieve similar winrates as REBEL, REBEL is more computationally tractable and simpler to implement.
5.2 General Chat
Task. We consider a general chat scenario where $x$ is a prompt from the user and $y$ is a response. To demonstrate the robustness of our method, we apply REBEL on two entirely different sets of models and datasets. For the first set, we adapt the setting from Zhu et al. (2023), using OpenChat-3.5 (Wang et al., 2024) as the base model, Starling-RM-7B-alpha (Zhu et al., 2023) as the reward model, and the Nectar dataset (Zhu et al., 2023). This set allows us to compare REBEL to APA (Zhu et al., 2023). We compare our trained model with Starling-LM-7B-alpha. For the second set, we use Meta-Llama-3-8B-Instruct (Meta, 2024) as the base model, FsfairX-LLaMA3-RM-v0.1 (Xiong et al., 2024) as the reward model, and the UltraFeedback dataset (Cui et al., 2023).
Evaluation. Following previous works, we use AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2023), and Open LLM Leaderboard (Beeching et al., 2023) as metrics. AlpacaEval 2.0 uses prompts from AlpacaFarm (Dubois et al., 2024) to compare model responses against a reference response generated by GPT-4-Turbo. We report the winrate over the reference responses. MT-Bench consists of 80 open-ended questions on various topics. Answers are scored directly by GPT-4. Open LLM Leaderboard consists of MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), Arc (Clark et al., 2018), Winogrande (Sakaguchi et al., 2019), TruthfulQA (Lin et al., 2022), and HellaSwag (Zellers et al., 2019). The prompts of the tasks consist of zero or few-shot samples.
Model MT-Bench AlpacaEval 2.0
1st Turn 2nd Turn Average LC Win Rate Win Rate
Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) 7.78 7.25 7.51 17.1 14.7
OpenChat-3.5 (Wang et al., 2024) 8.03 7.35 7.69 12.2 11.7
Starling-LM-7B-alpha (Zhu et al., 2023) 8.18 6.69 7.43 14.7 14.2
Meta-Llama-3-8B-Instruct (Meta, 2024) 8.31 7.89 8.10 22.9 22.6
GPT-3.5-Turbo (OpenAI, 2023) - - 7.94 18.1 9.6
Gemini Pro (Google, 2024) - - - 24.4 18.2
Claude 2 (Anthropic, 2023) - - 8.06 28.2 17.2
GPT-4 (OpenAI, 2023) - - 8.96 35.3 22.1
REBEL-OpenChat-3.5 $8.54_{(+0.51)}$ $7.58_{(+0.23)}$ $8.06_{(+0.37)}$ $17.3_{(+5.1)}$ $12.8_{(+1.1)}$
REBEL-Llama-3 $8.63_{(+0.32)}$ $7.69_{(-0.20)}$ $8.16_{(+0.06)}$ $30.1_{(+7.2)}$ $32.6_{(+10.0)}$
Table 2: Results on General Chat. The values in parentheses denote the improvement from the respective base models. The colors group the base model with the respective REBEL finetuned model.
Model MMLU GSM8K Arc Winogrande TruthfulQA HellaSwag Average
(5-shot) (5-shot) (25-shot) (5-shot) (0-shot) (10-shot)
OpenChat-3.5 (Wang et al., 2024) 63.6 68.5 64.9 80.6 47.3 84.7 68.3
Starling-LM-7B-alpha (Zhu et al., 2023) 63.4 68.0 64.9 81.1 47.3 84.8 68.2
Meta-Llama-3-8B-Instruct (Meta, 2024) 65.8 75.3 62.0 75.5 51.7 78.7 68.2
REBEL-OpenChat-3.5 $63.7_{(+0.1)}$ $68.8_{(+0.3)}$ $64.3_{(-0.6)}$ $80.4_{(-0.2)}$ $48.2_{(+0.9)}$ $85.0_{(+0.3)}$ $68.4_{(+0.1)}$
REBEL-Llama-3 $65.8_{(+0.0)}$ $75.6_{(+0.3)}$ $61.7_{(-0.3)}$ $75.8_{(+0.3)}$ $51.7_{(+0.0)}$ $78.8_{(+0.1)}$ $68.2_{(+0.0)}$
Table 3: Results on Open LLM Leaderboard. The values in parentheses denote the improvement from the respective base models. The colors group the base model with the respective REBEL finetuned model.
Quality Analysis. The results between models trained with REBEL and baseline methods are shown in Table 2 and 3. For MT-Bench and AlpacaEval 2.0, under the same setup, REBEL-OpenChat-3.5 outperforms Starling-LM-7B-alpha which is trained with APA (Zhu et al., 2023) on both metrics, demonstrating the effectiveness of REBEL under chat setting and its superior performance over APA. REBEL-Llama-3 outperforms several state-of-the-art models including GPT-3.5-Turbo, Genimi Pro, and Claude 2 with only 8B parameters. For the metrics on Open LLM Leaderboard, REBEL is able to consistently enhance the performance of GSM8K and HellaSwag under different settings and maintain the overall average as the base models. Similar values on MMLU as base models indicate we preserve the basic capability of the pre-trained model during the RL fine-tuning process. We include a breakdown of MT-Bench in Appendix K.
@@ 5.3 Image Generation
Task. We also consider the setting of image generation, where, given a consistency model (Song et al., 2023) and a target reward function, we seek to train the consistency model to output images that garner a higher reward. We use 45 common animals as generation prompts similar to Black et al. (2023); Oertell et al. (2024) and the latent consistency model (Luo et al., 2023) distillation of the Dreamshaper v7 model, a finetune of stable diffusion (Rombach et al., 2021).
Figure 3: Learning curves as a function of reward queries to the LAION aesthetic predictor. The colored areas represent 95% CIs.
@@
We compare REBEL to a clipped, policy gradient objective (Black et al., 2023; Fan et al., 2024; Oertell et al., 2024) with the aim to optimize aesthetic quality to obtain a high reward from the LAION aesthetic score predictor (Schuhmann, 2022). This baseline does not use critics or GAE for advantage estimates. However, the clipping objective is clearly motivated by PPO, and thus, we simply name this baseline as PPO.
@@
Evaluation. We evaluate on the reward under the LAION aesthetic reward model for an equal number of reward queries/samples generated and an equal number of gradient updates. The aesthetic predictor is trained to predict human-labeled scores of images on a scale of 1 to 10. Images that tend to have the highest reward are artwork. Following Agarwal et al. (2021), we report inter-quartile means (IQM) with 95% confidence intervals (CIs) across three seeds for both REBEL and PPO. The CIs were calculated with percentile bootstrap with stratified sampling over three random seeds.
@@
Figure 4: Generated images using PPO and REBEL during an intermediate checkpoint. At the same number of epochs, REBEL observes a higher reward under the reward model. This can further be seen by the more diverse background of images generated from REBEL with less training time.
@@
Quality Analysis. Figure 3 shows REBEL optimizes the consistency model faster during the beginning of training and eventually achieves a performance similar to that of PPO. For our experiments, we tuned both batch size and learning rate for our algorithms, testing batch sizes of $[4,8,16]$ per GPU and learning rates $[1\text{e}-4,3\text{e}-4,6\text{e}-4,1\text{e}-3]$. The main difference in implementation between PPO and REBEL is the replacement of the clipped PPO objective with our regression objective. To maximize LAION-predicted aesthetic quality, both REBEL and PPO transform a model that produces plain images into one that produces artistic drawings. We found across multiple seeds that REBEL produced lush backgrounds when compared to PPO’s generations. Please see Appendix G.3 for more examples of generated images.
@@ 6 Related Work
Policy Gradients. Policy gradient (PG) methods (Nemirovskij and Yudin, 1983; Williams, 1992; Konda and Tsitsiklis, 1999; Kakade, 2001; Schulman et al., 2017) are a prominent class of RL algorithms due to their direct, gradient-based policy optimization, robustness to model misspecification (Agarwal et al., 2020), and scalability to modern AI applications from fine-tuning LLMs (Stiennon et al., 2022) to optimizing text-to-image generators (Oertell et al., 2024).
Broadly speaking, we can taxonomize PG methods into two families. The first family is based on REINFORCE (Williams, 1992) and often includes variance reduction techniques (Kool et al., 2019; Richter et al., 2020; Zhu et al., 2023). While prior work by Ahmadian et al. (2024) has shown that REINFORCE-based approaches can outperform more complex RL algorithms like PPO on LLM fine-tuning tasks like TL;DR, we find that a properly optimized version of PPO still out-performs a REINFORCE baseline. The second family is adaptive PG techniques that precondition the policy gradient (usually with the inverse of the Fisher Information Matrix) to ensure it is covariant to re-parameterizations of the policy, which include NPG (Kakade, 2001; Bagnell and Schneider, 2003) and its practical approximations like TRPO (Schulman et al., 2015) and PPO (Schulman et al., 2017). Intuitively, the preconditioning ensures that we make small changes in terms of action distributions, rather than in terms of the actual policy parameters, leading to faster and more stable convergence. Unfortunately, computing and then inverting the Fisher Information Matrix is computationally intensive and therefore we often resort to approximations in practice, as done in TRPO. However, these approximations are still difficult to apply to large-scale generative models, necessitating even coarser approximations like PPO. In contrast, REBEL does not need any such approximations to be implemented at scale, giving us a much closer connection between theory and practice.
Reward Regression. The heart of REBEL is a novel reduction from RL to iterative squared loss regression. While using regression to fit either the reward (Peters and Schaal, 2007) or the value (Peng et al., 2019) targets which are then used to extract a policy have previously been explored, our method instead takes a page from DPO (Rafailov et al., 2023; Zhou et al., 2023) and inverse RL methods (Jacq et al., 2019; Watson et al., 2023) to implicitly parameterize the reward regressor in terms of the policy. This collapses the two-stage procedure of prior methods into a single step.
@@
The third family of PFT techniques combines elements from the previous two: it involves running an offline algorithm iteratively, collecting on-policy preference feedback from either a supervisor model such as GPT4 (Rosset et al., 2024; Xiong et al., 2024; Guo et al., 2024) or from a preference model fit on human data (Calandriello et al., 2024). All of these approaches can be considered instantiations of the general SPO reduction proposed by Swamy et al. (2024), which itself can be thought of as a preference-based variant of DAgger (Ross et al., 2011). Recent work by Tajwar et al. (2024) confirms the empirical strength of these techniques which leverage additional online data. Our approach fits best into this family of techniques – we also iteratively update our model by solving a sequence of supervised learning problems over on-policy datasets. However, REBEL comes with several key differentiating factors from the prior work. Online versions of DPO or IPO (Xiong et al., 2024; Tajwar et al., 2024; Guo et al., 2024; Calandriello et al., 2024; Munos et al., 2023) essentially use a reward / preference model to generate binary win-loss labels while REBEL actually uses the output of the reward model as a regression target, taking advantage of this more nuanced feedback. In contrast to Rosset et al. (2024), algorithmically, REBEL does not use any online preference feedback from GPT4 nor does it require to generate a large number of responses per prompt, both of which are extremely expensive as reported by Rosset et al. (2024). Theoretically, we are able to prove policy performance bounds under a much weaker coverage condition. Unlike Mao et al. (2024) that regularize to the initial policy $\pi_{0}$ during updates, we perform conservative updates by regularizing $\pi_{t+1}$ to $\pi_{t}$. When doing the former, it is difficult to prove convergence or monotonic improvement as the current policy can just bounce around a ball centered at $\pi_{0}$, a well-known issue in the theory of approximate policy iteration (Kakade and Langford, 2002; Munos, 2003). In contrast, by incorporating the prior policy’s probabilities into our regression problem, we are able to prove stronger guarantees for REBEL. When applying REBEL to the general preference setting, our algorithm shares some similarities with the one from Wu et al. (2024). The key difference is that by using paired samples, REBEL does not need to resort to approximations of the partition function (e.g., assume the partition function is always equal to a constant). Furthermore, we can run REBEL with datasets consisting of a mixture of on-policy and off-policy data with strong guarantees, enabling hybrid training, as previously explored in the RL (Song et al., 2023; Ball et al., 2023; Zhou et al., 2023) and inverse RL (Ren et al., 2024) literature.
7 Conclusion and Future Work